A circle has a diameter 27 cm. How far from the centre of this circle is a chord 18 cm long?
Answer by Cromlix(4381)

A circle with a diameter of 27cm
has a radius of 13.5cms
If we draw a line down to the centre
of the chord from the centre of the circle.
This will be AB.
Then we we connect a radius to the end of the
chord to the centre of the circle, we have a right angled
triangle with hypotenuse (radius) = 13.5cms.This
will be AC
A base side is half the size of the chord = 9cm.This
is BC
To find the distance from the centre of the circle
to the chord we must find the length of AB.
AB^2 + BC^2 = AC^2
AB^2 + 9^2 = 13.5^2
AB^2 = 13.5^2 - 9^2
AB^2 = 101.25
AB = square root of 101.25
AB = 10.1cms
This is the distance from the centre
of the circle to the chord.
Hope this helps.
:-)